OpenStudy (anonymous):

How long will it take to transfer 4.186*10^6 J of heat through 2 m^2 pane of 3mm thick glass(kglass=0.84) if temp difference is 10 degrees C? I can't get this one to solve correctly. I keep getting an option that is incorrect. Please help.

OpenStudy (anonymous):

\[Q [W] = -k \cdot A \cdot \left(\Delta T \over \Delta x \right)\]Divide the required amount of heat by the calculated Q to get the time.